help me solve the equation x+(x+2)+(x+4)= 54
Answer by checkley71(8403)

X+(X+2)+(X+4)=54
3X+6=54
3X=54-6
3X=48
X=48/3
X=16 ANSWER
PROOF:
16+(16+2)+(16+4)=54
16+18+20=54
54=54